the percentage uncertainity in mass and spring constant in time periodof spring mass respectively are 2% and 6 % .the percent uncertainity in its time period will be
(a) 3%
(b) 4%
(c) 8%
(d) none
1
Expert's answer
2019-07-01T09:22:16-0400
"T=2\pi \sqrt{\frac{m}{k}}"

"\frac{\Delta T}{T}=\frac{1}{2}\frac{\Delta m}{m}+\frac{1}{2}\frac{\Delta k}{k}"

"\frac{\Delta T}{T}=\frac{1}{2}0.02+\frac{1}{2}0.06=0.04=4\%"

Answer: (b) 4%